# Impact of the latest 22Ne+ reaction rates on nucleosynthesis in massive stars and galactic chemical evolution

### v1

@@ Abstract
In massive stars (initial mass of 9 ), the weak (slow neutron capture) process produces elements between Fe and Zr, enriching the Galaxy with these elements through core-collapse supernova explosions. The weak s-process nucleosynthesis is driven by neutrons produced in the 22Ne()25Mg reaction during convective He-core and C-shell burning. The yields of heavy elements thus depend on the 22Ne()25Mg and the competitive 22Ne()26Mg reaction rates which are dominated by several narrow-resonance reactions. While the accuracy of these rates has been under debate for decades, recent experimental efforts, including ours, drastically reduced these uncertainties. In this work, we use a set of 280 massive star nucleosynthesis models calculated using different 22Ne()25Mg and 22Ne()26Mg rates, and a galactic chemical evolution (GCE) study to probe their impact on the weak -process elemental abundances in the Galaxy. The GCE was computed with the OMEGA+ code, using the new sets of stellar yields with different 22Ne+ rates. From GCE, we find that these rates are causing up to 0.45 dex of variations in the [Cu/Fe], [Ga/Fe], and [Ge/Fe] ratios predicted at solar metallicity. The greatest impact on the stellar nucleosynthesis and GCE results derives from uncertainties in the (,) strength () of the Ex=11.32 MeV resonance. We show that the variations observed in the GCE calculations for the weak -proess elements become negligibly smaller than dispersions found in observations once the is accurately determined within the uncertaintiy of 10ā20% (typically reported experimental errors for the resonance) in future nuclear physics experiments.
@@ 1 Introduction
The slow neutron capture, namely the -process, produces about half of the solar abundances for elements heavier than Fe (Kappeler etĀ al., 2011; Lugaro etĀ al., 2023). The other half of the heavy elements are mostly produced by the rapid neutron capture process (r-process Woosley etĀ al., 1994; Kajino etĀ al., 2019; Cowan etĀ al., 2021; Siegel, 2022), with much smaller contributions from the p-process (Rauscher etĀ al., 2013; Pignatari etĀ al., 2016a). The contribution of neutrino-winds ejecta in core-collapse supernovae (CCSN) (Psaltis etĀ al., 2022, and references therein) and from the i-process (e.g., Denissenkov etĀ al., 2019; Choplin etĀ al., 2019) is still a matter of debate.
@@
The relevant stellar sources for the s-process elements are low- and intermediate-mass Asymptotic Giant Branch (AGB) stars and massive stars (initial mass 9 ). AGB stars produce the bulk of the solar s-process abundances between Sr and Bi (e.g., Bisterzo etĀ al., 2014; Prantzos etĀ al., 2020), while massive stars are responsible for most of the production between Fe and Sr (weak -process; e.g., Pignatari etĀ al., 2010). Most of the neutrons powering the s-process in AGB stars are generated by the 13CO reaction in the radiative 13C-pocket, with a smaller contribution by the 22NeMg reaction during Thermal Pulse events (e.g., Straniero etĀ al., 2006; Karakas etĀ al., 2014; Battino etĀ al., 2016; Lugaro etĀ al., 2023). In massive stars, 22NeMg is the main neutron source powering the s-process, which is efficiently activated in the convective He-burning core and the convective C-burning shell (e.g., Raiteri etĀ al., 1991b; Baraffe etĀ al., 1992; Kappeler etĀ al., 1994; Rauscher etĀ al., 2002; The etĀ al., 2007; Pignatari etĀ al., 2010). The 13CO reaction may be activated in the C-burning core, but its products are typically further processed by more advanced stages and not ejected by the CCSN explosion (Chieffi etĀ al., 1998; Pignatari etĀ al., 2013). Thus, the most relevant s-process stellar sources have been well-established. On the other hand, several uncertainties are affecting the stellar yields of s-process elements, mainly for CCSN. For these stars, quantitative predictions of theoretical stellar yields have been affected by uncertainties of both stellar models (e.g., Kappeler etĀ al., 1994; Pignatari etĀ al., 2016b; Sukhbold etĀ al., 2016; Limongi & Chieffi, 2018; Roberti etĀ al., 2024a) and nuclear reaction rates (Pignatari etĀ al., 2010; Nishimura etĀ al., 2017; Ota etĀ al., 2021; Wiescher etĀ al., 2023; Pignatari etĀ al., 2023, and references therein). The 22NeMg cross sections are particularly complicated because the reactions involve many resonances in their Gamow window. Therefore, their large uncertainties still have a significant impact on s-process nucleosynthesis (e.g., Wiescher etĀ al., 2012; Adsley etĀ al., 2021; Best etĀ al., 2025). Additionally, as the 22NeMg reaction is an endothermic reaction, the neutron emission channels only open above a certain energy (=478 keV, Wang etĀ al., 2021) where is a Center-of-Mass energy in 22Ne+ reaction). For resonances near the neutron threshold like this, -ray and neutron emission channels can compete. The n/ branching ratio of the resonance depends on the strength of its single-particle component (Wiescher etĀ al., 2023). 22Ne in the massive stars can be consumed by 22NeMg reactions, leading to less active 22NeMg reactions in the end (Kappeler etĀ al., 1994; Wiescher etĀ al., 2012). As such, 22NeMg cross sections also need to be constrained to understand the s-process yields from the CCSN.
@@
Like other charged-particle reactions, measuring these cross sections directly at stellar temperature is extremely challenging because of the small cross sections due to the extremely energy-dependent Coulomb barriers. The development of underground accelerator facilities, where neutron background environments are dramatically improved compared to those above the ground, is rapidly progressing (Best etĀ al., 2016; Gao & JUNA Collaboration, 2022; Shahina etĀ al., 2022). Nevertheless, the cross sections at low energy still need to rely on indirect measurements. Although both 22NeMg and 22NeMg cross sections have been studied by both direct (e.g., Wolke etĀ al., 1989; Harms etĀ al., 1991; Jaeger etĀ al., 2001; Hunt etĀ al., 2019; Shahina etĀ al., 2024) and indirect (e.g., Giesen etĀ al., 1993; Longland etĀ al., 2009; deBoer etĀ al., 2010; Massimi etĀ al., 2012; Talwar etĀ al., 2016; Massimi etĀ al., 2017; Adsley etĀ al., 2017, 2018; Jayatissa etĀ al., 2020; Ota etĀ al., 2020, 2021; Chen etĀ al., 2021) measurements for decades, many ambiguities remain.
@@
Unambiguously determining these rates will drastically reduce uncertainties of nucleosynthesis calculations in massive stars, which are essential for improving the robustness of CCSN yields. The impacts of different 22Ne+ reaction rates on stellar yields are frequently discussed, but limited to specific stellar conditions (e.g., Longland etĀ al., 2012; Talwar etĀ al., 2016; Adsley etĀ al., 2021; Ota etĀ al., 2021). Their direct impacts on galactic chemical evolution (GCE) have not yet been studied. In particular, we expect that the s-process in massive stars is the main producer of solar abundances of Cu, Ga, and Ge; its contributions to other elements are less certain due to various uncertainties, including those in the processing of s-process rich layers by the CCSN explosion (Pignatari etĀ al., 2010; Prantzos etĀ al., 2020).
@@
In the present work, we therefore perform simulations to study the GCE of Cu, Ga, and Ge and other elements (from Zn to Sr), particularly focusing on their sensitivity to the 22Ne + reaction rates in massive stars. Different 22Ne + rates, including a conventionally referenced rate (Longland etĀ al., 2012, hereafter LO12) and the rate we obtained from the experiment by Ota etĀ al. (2020) (hereafter OT20), are used in our simulations. Since many uncertainties remain in the reaction rates, multiple reaction rates that adopt different 22Ne + cross sections have been published (e.g., Adsley etĀ al., 2021; Wiescher etĀ al., 2023). OT20 produces the smallest weak s-process stellar yields among these rates, (Ota etĀ al., 2021) because the adopted rate for the () reaction is the smallest and that for the () reaction is relatively large. In contrast, the LO12 rate produces the largest weak -process yields among the four reaction rates mentioned here. Because the 22Ne + reaction in -process stars is dominated by resonance reactions, GCE sensitivity to 22Ne+ rates can be investigated by varying the strength of individual resonances. We estimate the impact of individual resonances on the GCE using this approach, comparing the GCE by LO12 with that by OT20. We aim to determine a desirable precision and upper limits for future measurements to constrain the strengths of the relevant 22Ne+ resonances. The same stellar set provided by Ritter etĀ al. (2018b) is used in our nucleosynthesis calculations for the GCE simulations. We thus focus on studying the impact of different reaction rates on the stellar abundances and GCE.
@@
Since the ground states of both 22Ne and have spin-parity () of 0+ and their reaction Q-value for populating the compound nucleus, 26Mg, is highly positive (Q = 10.614 MeV), 22Ne+ reaction primarily proceeds through resonant capture to natural-parity states in the highly excited 26Mg (see review by Wiescher etĀ al., 2023). While the compound nucleusās level density is high, only a limited number of resonances that have natural-parity and high -cluster strength are expected to contribute to the 22Ne+ reaction. The expectation has been verified by direct measurements, also confirming these resonancesā widths () are narrow and do not overlap between each resonance. For a given resonance, the key properties that determine its contribution to the stellar rate are the 22Ne+ resonance energy (Ex = + ECM) and the resonance strengths, and . Two (one possible) resonances, at =11.32 and 11.17 MeV, have been intensively studied due to their importance in the weak s-processās temperature (0.3ā0.35 GK). Although the weak s-process is reactivated at about 1 GK, the final abundances are less sensitive to the current uncertainties of the resonance strengths at the relevant energy.
@@
One of the two important resonances is well-identified at ECM = 0.702 MeV (Ex = 11.32 MeV). Both the () and () stellar rates are dominated by this resonance. It has been observed in both direct () (Harms etĀ al., 1991; Drotleff etĀ al., 1991; Giesen etĀ al., 1993; Jaeger etĀ al., 2001; Shahina etĀ al., 2024) and () experiments (Wolke etĀ al., 1989; Hunt etĀ al., 2019; Shahina etĀ al., 2022). Its is well established with these published measurements in excellent agreement with each other (Best etĀ al., 2025). In contrast, extracted from the direct () measurements are in poor agreement, making the for the Ex = 11.32 MeV resonance the outstanding contributor to uncertainties on the total 22Ne()25Mg stellar rate.
@@
At lower temperatures (0.25 GK), both reaction channels may be dominated by one or more resonances between the neutron threshold (=11.09 MeV) and ECM 0.635 MeV (Ex 11.25 MeV). In this region, in particular the resonance at Ex 11.17 MeV has long been controversial (see e.g., Longland etĀ al., 2012) since reported by Berman etĀ al. (1969). Jaeger etĀ al. (2001) searched for this state in their direct () measurement, setting an upper limit on the at Ex 11.17 MeV, 60 neV. Recently, Shahina etĀ al. (2022) also observed no strong transitions around this energy region in their () experiment, setting an upper limit of 150 neV on a candidate resonance at Ex 11.17 MeV. Additionally, the for the natural-parity resonances identified by Massimi etĀ al. (2012), the lowest-lying resonance above , = 11.11 MeV, remain poorly constrained (60 neV) by direct measurements (Jaeger etĀ al., 2001). Although this resonance has never been reported to have strong -cluster strength by indirect measurements, it may have the potential to completely dominate the () rate below 0.3 GK due to its proximity to the Gamow peak (Ota etĀ al., 2021). While the = 11.32 MeV and 11.17 MeV resonances are critical for 0.2ā0.35 GK, two resonances below the neutron threshold (therefore no () cross sections), 11.08 and 10.95 MeV resonances, could contribute to consume 22Ne via () reactions at 0.1ā0.2 GK (Ota etĀ al., 2020) before () reaction is activated at a later (hotter) stage of the star (Kappeler etĀ al., 1994; Wiescher etĀ al., 2023).
@@ 3.1 Stellar models and CCSN explosions
For the nucleosynthesis calculations presented in this work, we used the 20 massive star models from the stellar set by Ritter etĀ al. (2018b) 111Initial mass =12, 15, 20, and 25 (where ZAMS means Zero Age Main Sequence) and metallicity = 0.0001, 0.001, 0.006, 0.01, and 0.02 calculated with the MESA code, rev. 3372 (Paxton etĀ al., 2011). The CCSN models are calculated with a semi-analytical approach based on the Sedov blast wave (SBW) solution for the determination of the peak velocity of the shock throughout the stellar structure (Sedov, 1946), while the remnant mass is determined using the delayed explosion prescription from Fryer etĀ al. (2012). For further detail, readers are referred to Pignatari etĀ al. (2016b), Ritter etĀ al. (2018b) and Roberti etĀ al. (2024b).
For =0.01 and 0.02, the initial abundances adopted are solar-scaled (Grevesse & Noels, 1993). At = 0.006 and below, -enhanced abundances were used, where the isotopes of 12C, 16O, 20Ne, 24Mg, 28Si, 32S, 36Ar, 40Ca, and 48Ti are enhanced compared to a solar scaled composition (see Table 1, in Ritter etĀ al., 2018b), based on stellar observations. Different -enhancements are applied to different species, varying between [Ca/Fe]=+0.222 (where the enhancement is applied to 40Ca, and the remaining Ca isotopes are solar scaled), and [O/Fe]=+0.886 (the enhancement is only applied to 16O). Although the solar abundances by Grevesse & Noels (1993) have been updated several times (more recently, see Magg etĀ al., 2022), these variations would not affect the results of the present study.
The Multi-zone Post-Processing Nucleosynthesis in Parallel code (MPPNP) developed by the NuGrid collaboration is used for the calculations of the complete nucleosynthesis (e.g., Pignatari etĀ al., 2016b) as used in Ritter etĀ al. (2018b). The MPPNP post-processes the production and destruction of all relevant isotopes during stellar evolution and during CCSN explosions, allowing us to derive stellar isotopic yields. In this work, the nucleosynthesis simulations are performed for the aforementioned stellar models, where only the 22Ne+ rates are modified. The stellar structure and CCSN models can be used unchanged for different 22Ne+ rates, as the reactions do not impact energy generation and therefore do not affect stellar evolution properties. Both the pre-SN and post-SN nucleosynthesis were calculated using version V2.2 of the JINA Reaclib database if reaction rates from experiments were not available. For instance, for the 22Ne(,n)25Mg rate and the 22Ne(,)26Mg rate, Ritter etĀ al. (2018b) adopted Jaeger etĀ al. (2001) and Angulo etĀ al. (1999), respectively. Experimental neutron capture reaction rates were taken, when available, from the KADoNIS compilation (Dillmann etĀ al., 2006). We refer to Pignatari etĀ al. (2016b) and Ritter etĀ al. (2018b) for more details of all the reaction rates used for the post-processing calculations.

### v2

@@ Abstract
In massive stars (initial mass of 9 ), the weak (slow neutron capture) process produces elements between Fe and Zr, enriching the Galaxy with these elements through core-collapse supernova explosions. The weak s-process nucleosynthesis is driven by neutrons produced in the 22Ne()25Mg reaction during convective He-core and C-shell burning. The yields of heavy elements thus depend on the 22Ne()25Mg and the competitive 22Ne()26Mg reaction rates which are dominated by several narrow-resonance reactions. While the accuracy of these rates has been under debate for decades, recent experimental efforts, including ours, drastically reduced these uncertainties. In this work, we use a set of 280 massive star nucleosynthesis models calculated using different 22Ne()25Mg and 22Ne()26Mg rates, and a galactic chemical evolution (GCE) study to probe their impact on the weak -process elemental abundances in the Galaxy. The GCE was computed with the OMEGA+ code, using the new sets of stellar yields with different 22Ne+ rates. From GCE, we find that these rates are causing up to 0.45 dex of variations in the [Cu/Fe], [Ga/Fe], and [Ge/Fe] ratios predicted at solar metallicity. The greatest impact on the stellar nucleosynthesis and GCE results derives from uncertainties in the (,) strength () of the Ex=11.32 MeV resonance. We show that the variations observed in the GCE calculations for the weak -proess elements become negligibly smaller than dispersions found in observations once the is accurately determined within the uncertaintiy of 10–20% (typically reported experimental errors for the resonance) in future nuclear physics experiments.
@@ 1 Introduction
The slow neutron capture, namely the -process, produces about half of the solar abundances for elements heavier than Fe (Kappeler2011; lugaro:23). The other half of the heavy elements are mostly produced by the rapid neutron capture process (r-process Woosley1994; kajino:19; cowan:21; Siegel2022), with much smaller contributions from the p-process (rauscher:13; pignatari:16). The contribution of neutrino-winds ejecta in core-collapse supernovae (CCSN) (psaltis:22, and references therein) and from the i-process (e.g., Denissenkov2019; Choplin2019) is still a matter of debate.
@@
The relevant stellar sources for the s-process elements are low- and intermediate-mass Asymptotic Giant Branch (AGB) stars and massive stars (initial mass 9 ). AGB stars produce the bulk of the solar s-process abundances between Sr and Bi (e.g., bisterzo:14; Prantzos2020), while massive stars are responsible for most of the production between Fe and Sr (weak -process; e.g., 2010ApJ...710.1557P). Most of the neutrons powering the s-process in AGB stars are generated by the 13CO reaction in the radiative 13C-pocket, with a smaller contribution by the 22NeMg reaction during Thermal Pulse events (e.g., Straniero2006; karakas:14; battino:16; lugaro:23). In massive stars, 22NeMg is the main neutron source powering the s-process, which is efficiently activated in the convective He-burning core and the convective C-burning shell (e.g., raiteri:91; baraffe:92; Kappeler1994; rauscher:02; the:07; 2010ApJ...710.1557P). The 13CO reaction may be activated in the C-burning core, but its products are typically further processed by more advanced stages and not ejected by the CCSN explosion (chieffi:98; Pignatari2013). Thus, the most relevant s-process stellar sources have been well-established. On the other hand, several uncertainties are affecting the stellar yields of s-process elements, mainly for CCSN. For these stars, quantitative predictions of theoretical stellar yields have been affected by uncertainties of both stellar models (e.g., Kappeler1994; 2016ApJS..225...24P; sukhbold:16; limongi:18; roberti:24) and nuclear reaction rates (2010ApJ...710.1557P; nishimura:17; Ota2021; Wiescher2023; pignatari:23, and references therein). The 22NeMg cross sections are particularly complicated because the reactions involve many resonances in their Gamow window. Therefore, their large uncertainties still have a significant impact on s-process nucleosynthesis (e.g., Wiescher2012; Adsley2021; Best2025). Additionally, as the 22NeMg reaction is an endothermic reaction, the neutron emission channels only open above a certain energy (=478 keV, Wang2021) where is a Center-of-Mass energy in 22Ne+ reaction). For resonances near the neutron threshold like this, -ray and neutron emission channels can compete. The n/ branching ratio of the resonance depends on the strength of its single-particle component (Wiescher2023). 22Ne in the massive stars can be consumed by 22NeMg reactions, leading to less active 22NeMg reactions in the end (Kappeler1994; Wiescher2012). As such, 22NeMg cross sections also need to be constrained to understand the s-process yields from the CCSN.
@@
Like other charged-particle reactions, measuring these cross sections directly at stellar temperature is extremely challenging because of the small cross sections due to the extremely energy-dependent Coulomb barriers. The development of underground accelerator facilities, where neutron background environments are dramatically improved compared to those above the ground, is rapidly progressing (Best2016; Gao2022; Shahina2022). Nevertheless, the cross sections at low energy still need to rely on indirect measurements. Although both 22NeMg and 22NeMg cross sections have been studied by both direct (e.g., Wolke1989; Harms1991; Jaeger2001; Hunt2019; Shahina2024) and indirect (e.g., Giesen1993; Longland2009; deBoer2010; Massimi2012; Talwar2016; Massimi2017; Adsley2017; Adsley2018; Jayatissa2020; OTA2020135256; Ota2021; Chen2021) measurements for decades, many ambiguities remain.
@@
Unambiguously determining these rates will drastically reduce uncertainties of nucleosynthesis calculations in massive stars, which are essential for improving the robustness of CCSN yields. The impacts of different 22Ne+ reaction rates on stellar yields are frequently discussed, but limited to specific stellar conditions (e.g., Longland2012; Talwar2016; Adsley2021; Ota2021). Their direct impacts on galactic chemical evolution (GCE) have not yet been studied. In particular, we expect that the s-process in massive stars is the main producer of solar abundances of Cu, Ga, and Ge; its contributions to other elements are less certain due to various uncertainties, including those in the processing of s-process rich layers by the CCSN explosion (2010ApJ...710.1557P; Prantzos2020).
@@
In the present work, we therefore perform simulations to study the GCE of Cu, Ga, and Ge and other elements (from Zn to Sr), particularly focusing on their sensitivity to the 22Ne + reaction rates in massive stars. Different 22Ne + rates, including a conventionally referenced rate (Longland2012, hereafter LO12) and the rate we obtained from the experiment by OTA2020135256 (hereafter OT20), are used in our simulations. Since many uncertainties remain in the reaction rates, multiple reaction rates that adopt different 22Ne + cross sections have been published (e.g., Adsley2021; Wiescher2023). OT20 produces the smallest weak s-process stellar yields among these rates, (Ota2021) because the adopted rate for the () reaction is the smallest and that for the () reaction is relatively large. In contrast, the LO12 rate produces the largest weak -process yields among the four reaction rates mentioned here. Because the 22Ne + reaction in -process stars is dominated by resonance reactions, GCE sensitivity to 22Ne+ rates can be investigated by varying the strength of individual resonances. We estimate the impact of individual resonances on the GCE using this approach, comparing the GCE by LO12 with that by OT20. We aim to determine a desirable precision and upper limits for future measurements to constrain the strengths of the relevant 22Ne+ resonances. The same stellar set provided by Ritter2018 is used in our nucleosynthesis calculations for the GCE simulations. We thus focus on studying the impact of different reaction rates on the stellar abundances and GCE.
@@
Since the ground states of both 22Ne and have spin-parity () of 0+ and their reaction Q-value for populating the compound nucleus, 26Mg, is highly positive (Q = 10.614 MeV), 22Ne+ reaction primarily proceeds through resonant capture to natural-parity states in the highly excited 26Mg (see review by Wiescher2023). While the compound nucleus’s level density is high, only a limited number of resonances that have natural-parity and high -cluster strength are expected to contribute to the 22Ne+ reaction. The expectation has been verified by direct measurements, also confirming these resonances’ widths () are narrow and do not overlap between each resonance. For a given resonance, the key properties that determine its contribution to the stellar rate are the 22Ne+ resonance energy (Ex = + ECM) and the resonance strengths, and . Two (one possible) resonances, at =11.32 and 11.17 MeV, have been intensively studied due to their importance in the weak s-process’s temperature (0.3–0.35 GK). Although the weak s-process is reactivated at about 1 GK, the final abundances are less sensitive to the current uncertainties of the resonance strengths at the relevant energy.
@@
One of the two important resonances is well-identified at ECM = 0.702 MeV (Ex = 11.32 MeV). Both the () and () stellar rates are dominated by this resonance. It has been observed in both direct () (Harms1991; Drotleff1991; Giesen1993; Jaeger2001; Shahina2024) and () experiments (Wolke1989; Hunt2019; Shahina2022). Its is well established with these published measurements in excellent agreement with each other (Best2025). In contrast, extracted from the direct () measurements are in poor agreement, making the for the Ex = 11.32 MeV resonance the outstanding contributor to uncertainties on the total 22Ne()25Mg stellar rate.
@@
At lower temperatures (0.25 GK), both reaction channels may be dominated by one or more resonances between the neutron threshold (=11.09 MeV) and ECM 0.635 MeV (Ex 11.25 MeV). In this region, in particular the resonance at Ex 11.17 MeV has long been controversial (see e.g., Longland2012) since reported by Berman1969. Jaeger2001 searched for this state in their direct () measurement, setting an upper limit on the at Ex 11.17 MeV, 60 neV. Recently, Shahina2022 also observed no strong transitions around this energy region in their () experiment, setting an upper limit of 150 neV on a candidate resonance at Ex 11.17 MeV. Additionally, the for the natural-parity resonances identified by Massimi2012, the lowest-lying resonance above , = 11.11 MeV, remain poorly constrained (60 neV) by direct measurements (Jaeger2001). Although this resonance has never been reported to have strong -cluster strength by indirect measurements, it may have the potential to completely dominate the () rate below 0.3 GK due to its proximity to the Gamow peak (Ota2021). While the = 11.32 MeV and 11.17 MeV resonances are critical for 0.2–0.35 GK, two resonances below the neutron threshold (therefore no () cross sections), 11.08 and 10.95 MeV resonances, could contribute to consume 22Ne via () reactions at 0.1–0.2 GK (OTA2020135256) before () reaction is activated at a later (hotter) stage of the star (Kappeler1994; Wiescher2023).
These situations are well-described (in e.g., OTA2020135256; Adsley2021). In this work, we will therefore study the sensitivity of the GCE to the strengths of these individual resonances ( and ), including those above 11.32 MeV, by varying their strengths.
@@ 3.1 Stellar models and CCSN explosions
For the nucleosynthesis calculations presented in this work, we used the 20 massive star models from the stellar set by Ritter2018 111Initial mass =12, 15, 20, and 25 (where ZAMS means Zero Age Main Sequence) and metallicity = 0.0001, 0.001, 0.006, 0.01, and 0.02 calculated with the MESA code, rev. 3372 (2011ApJS..192....3P). The CCSN models are calculated with a semi-analytical approach based on the Sedov blast wave (SBW) solution for the determination of the peak velocity of the shock throughout the stellar structure (Sedov1946), while the remnant mass is determined using the delayed explosion prescription from Fryer2012. For further detail, readers are referred to 2016ApJS..225...24P, Ritter2018 and Roberti2024.
For =0.01 and 0.02, the initial abundances adopted are solar-scaled (Grevesse1993). At = 0.006 and below, -enhanced abundances were used, where the isotopes of 12C, 16O, 20Ne, 24Mg, 28Si, 32S, 36Ar, 40Ca, and 48Ti are enhanced compared to a solar scaled composition (see Table 1, in Ritter2018), based on stellar observations. Different -enhancements are applied to different species, varying between [Ca/Fe]=+0.222 (where the enhancement is applied to 40Ca, and the remaining Ca isotopes are solar scaled), and [O/Fe]=+0.886 (the enhancement is only applied to 16O). Although the solar abundances by Grevesse1993 have been updated several times (more recently, see Magg2022), these variations would not affect the results of the present study.
The Multi-zone Post-Processing Nucleosynthesis in Parallel code (MPPNP) developed by the NuGrid collaboration is used for the calculations of the complete nucleosynthesis (e.g., 2016ApJS..225...24P) as used in Ritter2018. The MPPNP post-processes the production and destruction of all relevant isotopes during stellar evolution and during CCSN explosions, allowing us to derive stellar isotopic yields. In this work, the nucleosynthesis simulations are performed for the aforementioned stellar models, where only the 22Ne+ rates are modified. The stellar structure and CCSN models can be used unchanged for different 22Ne+ rates, as the reactions do not impact energy generation and therefore do not affect stellar evolution properties. Both the pre-SN and post-SN nucleosynthesis were calculated using version V2.2 of the JINA Reaclib database if reaction rates from experiments were not available. For instance, for the 22Ne(,n)25Mg rate and the 22Ne(,)26Mg rate, Ritter2018 adopted Jaeger2001 and Angulo1999, respectively. Experimental neutron capture reaction rates were taken, when available, from the KADoNIS compilation (Dillmann2006). We refer to 2016ApJS..225...24P and Ritter2018 for more details of all the reaction rates used for the post-processing calculations.